An invertible square matrix A satisfies A^3 +3A^2 −25A+21I = O, where I and O are the identity and zero matrices, respectively.

Find the inverse of A^2
Answer #1

we have given. A3 + 3A? 25 A + 21 I = 0 A3 + 3 A2-25 A = -215 muuriply A-l both sidee. A? +3 A -251 = -21A-! agan muereply by